OpenStudy (lilshane):

Find both unit rates. 4543.08 km in 52.4 h

OpenStudy (lilshane):

@TheSmartOne

OpenStudy (lilshane):

@Michele_Laino

OpenStudy (lilshane):

@KendrickLamar2014

TheSmartOne (thesmartone):

4543.08 km in 52.4 h \(\sf\Large \frac{4543.08}{4543.08}~ km ~in ~\frac{52.4}{4543.08} ~h\) and \(\sf\Large \frac{4543.08}{52.4}~ km ~in ~\frac{52.4}{52.4} ~h\)

OpenStudy (lilshane):

I dont have that as an answer

OpenStudy (lilshane):

@TheSmartOne

TheSmartOne (thesmartone):

Simplify what I gave you :)

OpenStudy (lilshane):

I cant

TheSmartOne (thesmartone):

Use a calculator, divide up the fractions...